Solve the problem with complete solution. A brass wire is .500 m long at 43 degree Celsius. If heated at 70 degree Celsius, what is the change in length?


1
Expert's answer
2021-12-28T08:44:32-0500

Given

Initial length: "L_0=0.5\text{ m}."

Initial temperature: "t_1=43°\text C."

Final temperature: "t_2=70°\text C."

Coefficient of thermal expansion for brass: "\alpha=18·10^{-6}/°\text C."


Solution

Write the equation for thermal expansivity:

"\Delta L=L_0\alpha \Delta t=L_0\alpha(t_2-t_1),\\
\Delta L=0.5(18·10^{-6})(70-43)=2.43·10^{-4}\text{ m}."


Answer

0.243 mm.